OpenStudy (anonymous):

What is the value of 5 1/7 multiplied by 2 3/9 in simplest form?

OpenStudy (mertsj):

\[\frac{36}{7}\times\frac{21}{9}=\]

OpenStudy (mertsj):

Cancel out the common factors and multiply.

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (mertsj):

yw